OpenStudy (anonymous):

NEEED HELP!! Using the quadratic formula to solve x^2 + 20 = 2x, what are the values of x?

OpenStudy (jdoe0001):

\(\bf x^2 + 20 = 2x\implies x^2-2x+20=0 \\ \quad \\ \textit{quadratic formula}\\ {\color{blue}{ 1}}x^2{\color{red}{ -2}}x{\color{green}{ +20}}=0 \qquad \qquad x= \cfrac{ - {\color{red}{ b}} \pm \sqrt { {\color{red}{ b}}^2 -4{\color{blue}{ a}}{\color{green}{ c}}}}{2{\color{blue}{ a}}}\)

OpenStudy (anonymous):

i plug it in??